### v1

# DeepSport: A Multimodal Large Language Model for Comprehensive Sports Video Reasoning via Agentic Reinforcement Learning

Abstract
Sports video understanding presents unique challenges, requiring models to perceive high-speed dynamics, comprehend complex rules, and reason over long temporal contexts. While Multimodal Large Language Models (MLLMs) have shown promise in genral domains, the current state of research in sports remains narrowly focused: existing approaches are either single-sport centric, limited to specific tasks, or rely on training-free paradigms that lack robust, learned reasoning process. To address this gap, we introduce DeepSport, the first end-to-end trained MLLM framework designed for multi-task, multi-sport video understanding. DeepSport shifts the paradigm from passive frame processing to active, iterative reasoning, empowering the model to “think with videos” by dynamically interrogating content via a specialized frame-extraction tool. To enable this, we propose a data distillation pipeline that synthesizes high-quality Chain-of-Thought (CoT) trajectories from 10 diverse data source, creating a unified resource of 78k training data. We then employ a two-stage training strategy, Supervised Fine-Tuning (SFT) followed by Reinforcement Learning (RL) with a novel gated tool-use reward, to optimize the model’s reasoning process. Extensive experiments on the testing benchmark of 6.7k questions demonstrate that DeepSport achieves state-of-the-art performance, significantly outperforming baselines of both proprietary model and open-source models. Our work establishes a new foundation for domain-specific video reasoning to address the complexities of diverse sports.
footnotetext: Equal contribution.
1 Introduction
The domain of sports analysis has long served as a fertile ground for interdisciplinary research, driving advancements in both Computer Vision (CV) and Natural Language Processing (NLP) [THOMAS20173]. Traditional CV approaches have achieved remarkable success in specific, isolated tasks, such as fine-grained action detection and recognition [shao2020finegym, giancola2018soccernet], event spotting [Xarles_2024_CVPR], highlight generation [merler2018excitement, merler2018automatic], and player and ball trajectory detection and tracking in team sports [cioppa2021cameracalibrationplayerlocalization, vandeghen2022semi, 10411638, cite-key]. Simultaneously, NLP techniques have been effectively deployed for automated match reporting [huang2020generating] and sentiment analysis [vujivcic2023approach]. Recently, the emergence of Multimodal Large Language Models (MLLMs) [openai2024gpt4o, geminiteam2024gemini] has accelerated this trend, enabling more complex applications in sports [xia2024language], such as AI-assisted refereeing [held2023vars, held2023vars] and commentary generation [rao2024matchtimeautomaticsoccergame, you2025timesoccer].
In response to these capabilities, the community has produced a surge of high-quality, public benchmarks covering a wide array of disciplines. These range from action quality assessment in diving [xu2022finediving] to tactical classification in fencing [lai2024facts], and comprehensive question answering in large-scale benchmarks [xia2024sportqa, xia2024sportu, rao2025multi]. These datasets provide the necessary foundation for evaluating sophisticated sports intelligence.
However, despite this proliferation of data and benchmarks, the development of capable models has lagged behind. The current sports-specific MLLMs exhibits a significant fragmentation. For example, trained models are predominantly soccer-centric [rao2025multi, you2025timesoccer], optimizing for specific tasks like commentary generation within a single sport, but failing to generalize to others. On the other hand, attempts at multi-sport reasoning, such as FineQuest [chen2025finequestadaptiveknowledgeassistedsports], rely on training-free paradigms. Consequently, to the best of our knowledge, a unified, end-to-end trained MLLM capable of performing multiple tasks (e.g., commentary, foul detection, QA) across diverse sports does not yet exist.
Figure 1: Overview of the DeepSport framework. Trained through SFT and RL, DeepSport decides whether to utilize tools, supporting (a) passive single-pass inference, (b) single tool-use for two-turn interactions, and (c) iterative multi-turn reasoning via multiple tool calls. As shown on the Left, we categorize tasks into four core dimensions: Fine-Grained Recognition, Rule & Procedural Logic, Assessment & Coaching, and Live Commentary, covering diverse fine-grained sub-tasks (in red) across multiple sports.
To bridge this gap, we introduce DeepSport, the first sport-specific MLLM framework explicitly trained for multi-task, multi-sport video understanding (Figure 1). Inspired by the emerging paradigm of “Thinking with Videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong, ge2025framemindframeinterleavedvideoreasoning], DeepSport shifts from passive frame processing to active, iterative video interrogation via a specialized frame-extraction tool. To enable this, we devise a data distillation pipeline that synthesizes high-quality Chain-of-Thought (CoT) trajectories from 10 diverse data sources (derived from 9 existing works) covering 12 different sports, establishing a unified resource for training. We then employ a two-stage strategy—Supervised Fine-Tuning (SFT) followed by Agentic Reinforcement Learning with a novel gated reward function—to teach the model how to reason and when to use tools effectively.
Our main contributions are summarized as follows:
  • The First Multi-Task, Multi-Sport Trained MLLM. We propose DeepSport, a unified model framework that generalizes across diverse sports and four core capabilities: Fine-Grained Recognition, Rule & Procedural Knowledge, Assessment & Coaching, and Live Commentary & Reporting. It represents a paradigm shift from static, single-sport models to dynamic, tool-augmented sports intelligence.
  • A New Training Dataset and Evaluation Benchmark. We introduce a pipeline to curate, unify, and distill data from 10 existing sources, covering 12 different sports, including American football, badminton, baseball, basketball, boxing, diving, fencing, ice hockey, soccer, table tennis, volleyball, and gymnastics. In total, we construct 78K training data and 6.7K testing data.
  • Tool-Based Reinforcement Learning Strategy. We design a specialized reinforcement learning framework using Group Relative Policy Optimization (GRPO). We introduce a gated tool-use reward function that incentivizes the model to perform active, multi-turn visual reasoning while maintaining logical consistency.
  • State-of-the-Art Performance. Our DeepSport achieves an average score of 38.29, significantly outperforming all baselines, including general state-of-the-art (SOTA) models, and establishes a new SOTA for sports video tasks.
2 Related Work
Analyzing sports is fundamentally a dynamic, temporal reasoning problem. Core tasks, such as foul detection in basketball (e.g., traveling) or technical scoring in diving and gymnastics, are defined by sequences of motion over time, not by isolated static image. Consequently, sports analysis necessitates models that can parse and reason over complex video streams.
In this domain, early sports video analysis relied on task-specific, traditional computer vision (CV) pipelines. For example, for player detection and tracking [cioppa2021cameracalibrationplayerlocalization, liu2023automatedplayeridentificationindexing, cioppa2022soccernet], rally prediction [xia2022vren, 10411638], automated commentary generation[huang2020generating, 10871604, 10.1145/3341105.3374063], fine-grained action recognition in specific sports [shao2020finegym, wu2022survey, WANG2023e18124, giancola2018soccernet]. While effective for their spesific tasks, these specialized models lack unified reasoning capabilities and cross-modal flexibility.
2.1 Video Understanding Using Multimodal Large Language Models
Early efforts in general video understanding extended MLLMs to handle video inputs, enabling video conversation and description (e.g., Video-ChatGPT [Maaz2023VideoChatGPT], Video-LLaVA [lin2023video]). However, these open source models usually fall behind SOTA commercial MLLMs, such as GPT [gpt4o], Gemini[geminiteam2024gemini], and Claude [Claude3.5] in most video tasks, especially those that require complicated reasoning on multiple video question answer benchmarks [fang2024mmbenchvideolongformmultishotbenchmark, zhou2025mlvubenchmarkingmultitasklong]. Recent developments have focused on combining reinforcement learning (RL) with MLLMs, pushing models to achieve stronger reasoning abilities and demonstrate enhanced performance on various video understanding tasks, such as Video-R1[feng2025video] and DeepVideo-R1 [park2025deepvideor1videoreinforcementfinetuning]. Inspired by the concept of “Thinking with images” proposed by OpenAI [a2025_thinking], recent works have begun using agentic reinforcement learning to train MLLM as an agent that can use tools to dynamically and iteratively interrogate visual content as part of its reasoning process [zhang2025landscapeagenticreinforcementlearning, fan2025grit, zheng2025deepeyesincentivizingthinkingimages]. This agentic concept has been extended to the temporal domain to handle long videos. Recent works introduce MLLMs that can “think with long videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong]. These models act as agents that can densely sample new video frames on demand or perform multi-turn frame spotlighting by invoking a visual toolbox. This allows the model to actively “re-watch” segments and dynamically query the video for relevant frames, drastically reducing the number of frames processed and mitigating hallucination.
While these general-domain video MLLMs demonstrate powerful reasoning on benchmarks, such as Video-Holmes [cheng2025videoholmesmllmthinklike], they are trained for general-purpose rather than domain-specific tasks. The current general-domain MLLMs have been shown lacking the specialized domain knowledge required to understand the intricate rules, fine-grained actions, and professional terminology of diverse sports [xia2024sportu, rao2025multi].
2.2 Multimodal Large Language Models in Sports
Existing MLLMs in sports are typically limited in one of three ways: they are (i) training-free, relying on static knowledge graphs with limited tasks, such as no commentary generation task, which is one of the important things for sports viewers [chen2025finequestadaptiveknowledgeassistedsports]; (ii) single-sport, focusing on individual sports [rao2025multi, you2025timesoccer, kodathala2025sv33bsportsvideounderstanding, bao2025tennistvmultimodallargelanguage]; or (iii) single-task, such as models designed only for commentary [you2025timesoccer].
To the best of our knowledge, a single, end-to-end trained MLLM capable of performing multiple tasks across multiple sports does not yet exist. Therefore, our work introduces the first MLLM to address this challenge, leveraging tool-based agentic RL during training to unify these diverse capabilities.
Figure 2: DeepSport training overview. Given sport videos, we first perform data distillation with a teacher MLLM to construct DeepSport-CoT data and Supervised Fine-Tune a tool-augmented student model. We then further optimize the model with GRPO-based agentic reinforcement learning, where the agent iteratively calls a frame-extraction tool, produces chain-of-thought reasoning over new frames, and is guided by a reward manager that combines semantic accuracy, behavioral shaping, and format gating.
3 The DeepSport Framework
To address the gap in multi-task, multi-sport video understanding, we introduce DeepSport, an end-to-end trained MLLM trained with agentic reinforcement learning with a specific tool function, that extracts related frames and performs multi-turn generation. Our framework is built upon a two-stage training pipeline: (1) a Supervised Fine-Tuning (SFT) “cold-start” phase, followed by (2) an Agentic Reinforcement Learning (RL) phase. This SFT+RL paradigm is a validated approach for training capable “Thinking with Video” agents, as seen in recent works like FrameThinker [he2025framethinkerlearningthinklong]. Our methodology makes two primary contributions. First, we propose a data distillation pipeline to create a high-quality, large-scale dataset that includes the complex Chain-of-Thought (CoT) reasoning and tool-call syntax necessary for training the model. Second, we introduce a novel tool-based reward function within the Group Relative Policy Optimization (GRPO) framework to specifically incentivize effective multi-turn reasoning.
In this section, we first define our core paradigm (Section 3.1). We then detail our data distillation pipeline (Section 3.2) and conclude with the two-stage training strategy (Section 3.3).
3.1 Deepsport Video Reasoning Paradigm
Our framework empowers the MLLM to “think with videos” [zhang2025thinkingvideosmultimodaltoolaugmented, he2025framethinkerlearningthinklong]. The framework engages in a multi-turn reasoning loop, actively and strategically seeking information from the video for its decision-making.
The task can be defined as, given a user query $Q$ and a long video $V$, the model’s reasoning process generates a trajectory $\tau$. The process is initialized at $t=1$ by providing the model with an initial context $F_{1}$, which is a uniformly-sampled set of $k$ frames (e.g., $k=8$) from the entire video $V$. Each frame is tagged with its frame_index.
At each subsequent step $t$, the model must choose an action $A_{t}$ from its available action space $\mathcal{A}$, which consists of two types:
  • Frame Extraction Tool: <tool_call>
    frame_extraction_tool( $idx_{start},idx_{end}$)</tool_call>. This is an iterative action that allows the model to request a new set of frames from a specified temporal window.
  • Output Answer: <answer>...</answer>. This is a terminal action that concludes the reasoning trajectory and provides the final answer.
A complete trajectory $\tau$ is thus a sequence of $n$ steps, where each step consists of the frames provided to the model, the thought it generated, and the action it took:
$\tau=\big((F_{1},T_{1},A_{1}),(F_{2},T_{2},A_{2}),\dots,(F_{n},T_{n},A_{n})\big).$ (1)
Each triplet $(F_{t},T_{t},A_{t})$ at step $t$ is defined as:
  • Provided Frames ($F_{t}$): The set of $k$ frames available at the beginning of step $t$. For $t=1$, this is the initial sparse context ($F_{1}$). For $t>1$, this is the set of frames $F_{t}$ returned by the tool call $A_{t-1}$.
  • Thought ($T_{t}$): The model’s textual reasoning, generated within <think>...</think> tags. The model generates this thought based on $F_{t}$ and all prior history.
  • Action ($A_{t}$): The action chosen by the model after its thought $T_{t}$. If $A_{t}$ is the frame_extraction_tool, it generates $F_{t+1}$ for the next step. If $A_{t}$ is the output_answer, the trajectory terminates.
Inspired by the Cognitive Consistency Verification module in FrameThinker [he2025framethinkerlearningthinklong], we apply the following method to prevent of redundant exploration. If the model executes an action $A_{t}$ with the exact same temporal interval as its preceding action $A_{t-1}$, the framework treats this as an invalid “format error,” which is penalized during training.
3.2 Data Distillation Pipeline
A significant bottleneck in training a sport MLLMs for iterative reasoning is the lack of high-quality, large-scale sports-specific SFT data that demonstrates the desired step-by-step reasoning and tool-use behavior.
1. Data Generation. We first curated a diverse set of 9 existing works (Table 1) with 10 different data sources to cover a wide spectrum of 12 distinct sports, spanning team sports (Soccer, Basketball, Volleyball, American Football, Ice Hockey, Baseball), racket games (Table Tennis, Badminton), combat sports (Fencing, Boxing), and artistic disciplines (Diving, Gymnastics). We unified these disparate data sources to establish a comprehensive taxonomy of tasks across four high-level dimensions: (1) Fine-Grained Recognition: encompassing foundational visual perception tasks such as Action Sequencing, Player Identification, Camera Transition Analysis, etc.; (2) Rule & Procedural Logic: focusing on regulatory knowledge, including Foul Classification, Referee Decision Analysis, Penalty Description, etc.; (3) Assessment & Coaching: requiring expert-level reasoning for Tactical Analysis, Technical Error Identification, Scoring Prediction, etc.; and (4) Live Commentary & Reporting: covering multimodal narration tasks like Play-by-Play Commentary, Game Situation Analysis, Score Reading, etc.
Since several of these benchmarks are not in a question-answering format (FineDiving [xu2022finediving] for action quality, SoccerReplay-1988 [rao2024towards] for commentary generation, FACTS [lai2024facts] for action classification, or T3Set [10.1145/3711896.3737407] for tactic correction), our first step was to unify them. We designed a series of task-specific templates to convert their original labels into structured QA pairs.
Table 1: Source datasets curated for our data distillation pipeline. Tasks marked with (*) were converted into a QA format using templates.
Dataset Source Sport(s) Original Task / Content
SoccerBench [rao2025multi] Soccer Multi-task QA
SoccerReplay-1988 ([rao2024towards]) Soccer Commentary Generation (*)
SPORTU [xia2024sportu] Multi-sport Multi-level & -task QA
Sports-QA [li2024sports] Multi-sport Multi-task QA
SportR [xia2025sportrbenchmarkmultimodallarge] Multi-sport Foul & Tactic Reasoning QA
FineDiving [xu2022finediving] Diving Action Quality Assessment (*)
FACTS [lai2024facts] Fencing, Boxing Action Classification (*)
T3-Set ([10.1145/3711896.3737407]) Table Tennis Tactic Correction (*)
X-VARS([held2024x]) Soccer Foul Recognition & Explanation
2. Data Splitting and Integrity. We follow video-level splitting: all questions associated with a single video (even if there are multiple) are assigned exclusively to one data split. This prevents the model from being trained and tested on frames from the same video. Furthermore, we explicitly handle dataset overlap; for instance, any video clip used in the SoccerBench [rao2025multi] test set is excluded from our SoccerReplay-1988 [rao2024towards] training pool. The detailed number of QA split will be illustrated in Section 4.1.
3. CoT Data Distillation. With a unified set of QA prompts, we generate high-quality reasoning trajectories. We prompt the current SOTA model, Qwen3-VL-235B-A22B-Thinking [qwen3technicalreport], to generate detailed, step-by-step reasoning. Critically, the prompt instructs the model to invoke the frame_extraction_tool (defined in Section 3.1) when necessary. This process results in CoT trajectories that are interleaved with tool calls.
We employ an LLM-as-Judge (using DeepSeek-V3.2-Exp [deepseekai2024deepseekv32]) to filter for high-quality trajectories. The judge scores each generated CoT on the predicted answer. Only trajectories that pass a high-quality threshold are retained for SFT Training. As a result, we constructed the DeepSport-CoT-15k dataset, which contains 15,000 high-quality Q&A pairs with CoT annotations, to be used later in SFT training.
3.3 Two-Stage Training Strategy
We first perform Supervised Fine-Tuning (SFT) to teach the model the CoT format as the “cold start”, which has been wildly adopted in recent works [deepseekai2025deepseekr1incentivizingreasoningcapability, feng2025video, zhang2025thinkingvideosmultimodaltoolaugmented], followed by a Reinforcement Learning (RL) phase to optimize the model’s reasoning and tool-use policy.
3.3.1 SFT Cold-Start
We first perform an SFT phase using the DeepSport-CoT-15k. The objective of this phase is not to teach the model to solve all tasks, but to serve as a “cold-start” to familiarize the model with the complex CoT trajectory structure. By training on the DeepSport-CoT-15k, the model learns the syntax of our thought and action tags (e.g., <think>... and <tool_call>...) and the basic patterns of multi-turn reasoning. This initial alignment is crucial for stabilizing the subsequent RL phase, as it ensures the model starts with a reasonable policy capable of producing valid, well-formatted actions to explore.
3.3.2 Agentic RL with Gated Tool Rewards
In the second stage, we optimize the model’s policy using Reinforcement Learning.
Group Relative Policy Optimization (GRPO).
During the reinforcement learning phase, we employ the Group Relative Policy Optimization (GRPO) [shao2024deepseekmath], which has been commonly used in reinforcement learning for model training across multiple domains [deepseekai2025deepseekr1incentivizingreasoningcapability, liu2025finr1largelanguagemodel, feng2025video]. For each prompt $q$ from the RL set, we sample a group of $G$ trajectories $\{\tau_{i}\}_{i=1}^{G}\sim\pi_{\text{old}}(\cdot\mid q)$ and obtain rewards $\{r_{i}\}_{i=1}^{G}$. The advantage $A_{i}$ is defined by the group statistics:
$A_{i}\;=\;\frac{r_{i}-\mu_{r}}{\sigma_{r}}\!,$ (2)
where $\mu_{r}$ and $\sigma_{r}$ are the mean and standard deviation of the rewards $\{r_{i}\}$. Letting $\rho_{i}=\frac{\pi_{\theta}(\tau_{i}\mid q)}{\pi_{\text{old}}(\tau_{i}\mid q)}$, the GRPO objective is:
$\begin{split}\mathcal{J}_{\text{GRPO}}(\theta)=\mathbb{E}\Bigg[\frac{1}{G}\sum_{i=1}^{G}&\min\Big(\rho_{i}A_{i},\operatorname{clip}(\rho_{i},1-\varepsilon,1+\varepsilon)A_{i}\Big)\\ &-\beta\cdot\text{KL}\big(\pi_{\theta}(\cdot\mid q)\|\pi_{\text{ref}}(\cdot\mid q)\big)\Bigg].\end{split}$ (3)
where $\pi_{\text{ref}}$ is the reference model, which is usually the initial SFT model.
Gated Tool-Use Reward Function.
To train our model, we design a novel, multi-component reward function $R(\tau)$ that provides granular behavioral shaping to encourage both task success and effective reasoning.
1. Core Objective. The primary reward signal $R_{\text{acc}}$ measures task success, based on the semantic accuracy of the model’s final answer in trajectory $\tau$:
$R_{\text{acc}}(\tau)=acc(\tau),\quad acc(\tau)\in[0,1].$ (4)
Table 2: Performance comparison on the test set. We evaluate models across four high-level dimensions designed to assess holistic sports tasks. Each dimension encompasses diverse specific capabilities: Fine-Grained Recognition involves visual perception tasks (e.g., Action Recognition, Player Identification); Rule & Procedural tests regulatory knowledge (e.g., Foul Classification, Referee Decision Analysis); Assessment & Coaching evaluates expert reasoning (e.g., Tactical Analysis, Scoring Prediction); and Live Commentary & Reporting assesses multimodal narration (e.g., Play-by-Play Commentary, Score Reading).
Model Frames Fine-Grained Rule & Assessment & Live Commentary Overall
Recognition Procedural Coaching & Reporting Avg.
GPT-5 16 46.50 32.89 27.01 22.76 35.70
Qwen3-VL-8B-Instruct 16 43.82 23.49 25.52 20.77 31.66
Qwen3-VL-8B-Thinking 16 46.75 31.98 23.74 21.96 34.83
Video-R1 16 38.21 17.76 13.73 10.16 24.06
InternVL3.5-14B 16 46.38 29.10 22.64 17.59 33.01
Qwen3-VL-235B-A22B-Thinking 16 44.96 31.03 28.72 25.59 35.36
Qwen2.5-VL-7B-Instruct 16 28.20 13.71 9.09 3.13 16.98
DeepSport (Ours) 14.39 51.09 43.82 24.74 24.60 40.08
2. Behavioral Shaping. We introduce shaping rewards controlled by gating variables. We define the tool usage gate $g_{\text{tool}}(\tau)\in\{0,1\}$, which activates ($g_{\text{tool}}=1$) if the model invokes the frame_extraction_tool at least once. The accuracy gate $g_{\text{acc}}(\tau)$ activates if the accuracy is above a minimum threshold (e.g., $acc(\tau)\geq 0.5$):
$g_{\text{acc}}(\tau)=\begin{cases}1,&\text{if }acc(\tau)\geq 0.5,\\ 0,&\text{otherwise.}\end{cases}$ (5)
These gates define the tool usage incentive $R_{\text{tool}}$, which is designed to (a) strongly reward successful tool use, (b) provide a small “curiosity” bonus for exploratory tool use, and (c) give no reward for failing to use tools:
$R_{\text{tool}}(\tau)=\begin{cases}0.5\cdot acc(\tau),&\text{if }g_{\text{tool}}(\tau)=1\text{ and }g_{\text{acc}}(\tau)=1,\\[4.0pt] 0.03,&\text{if }g_{\text{tool}}(\tau)=1\text{ and }g_{\text{acc}}(\tau)=0,\\[4.0pt] 0,&\text{if }g_{\text{tool}}(\tau)=0.\end{cases}$ (6)
3. Format Gating and Final Reward. Finally, we enforce the structural and logical integrity of the trajectory (defined in Section 3.1). We use a format gate $g_{\text{fmt}}(\tau)\in\{0,1\}$ that validates the <tool_call>...</tool_call> syntax and penalizes invalid actions, such as redundant tool calls on the exact same interval. Invalid trajectories incur a structural penalty:
$P_{\text{format}}(\tau)=-0.05\,(1-g_{\text{fmt}}(\tau)).$ (7)
The total reward is then expressed as:
$R(\tau)=g_{\text{fmt}}(\tau)\big(R_{\text{acc}}(\tau)+R_{\text{tool}}(\tau)\big)+P_{\text{format}}(\tau).$ (8)
When the output format is valid ($g_{\text{fmt}}=1$), the reward combines task accuracy and tool-use behavior. Otherwise ($g_{\text{fmt}}=0$), the reward collapses to the penalty $R(\tau)=-0.05$.
4 Experiments
4.1 Experimental Setup
Datasets.
As described in Section 3.2, our data pipeline yields three distinct subsets based on a strict video-level split. First, we construct DeepSport-CoT-15k, consisting of 15k high-quality, LLM-filtered CoT QA pairs used specifically for the SFT cold-start phase. Second, for the reinforcement learning phase, we utilize DeepSport-RL-63k, which comprises 63k original, unified QA pairs (prompt plus ground truth answer) serving as the environment prompts. Finally, we evaluate performance on our held-out Testing Benchmark, containing 6.7k QA pairs that cover all tasks and sports to ensure a rigorous evaluation.
Implementation Details.
We select Qwen2.5-VL-7B [bai2025qwen2] as our backbone model, processing video inputs at a resolution of $640\times 360$. Our training pipeline consists of two stages: we first fine-tune the model on the DeepSport-CoT-15k set for 1 epoch, followed by reinforcement learning via our tool-based GRPO algorithm with a batch size of 32 and a rollout number of 8 on the DeepSport-RL-63k set for 300 steps. The model is trained to generate up to 12,800 tokens per response. Training was performed on 8 $\times$ H20 GPUs for totaling 796 GPU hours.
Baselines.
We compare DeepSport against a comprehensive suite of MLLMs, including both open- and closed-source models. Specifically, we select models from the Qwen family (Qwen2.5-VL-7B-Instruct as our primary baseline for direct comparison, Qwen3-VL-8B-Instruct/Thinking, and Qwen3-VL-235B-A22B-Thinking [bai2025qwen2, qwen3technicalreport]), as well as InternVL3.5-14B [wang2025internvl3_5] and Video-R1 [feng2025video]. For evaluation, we uniformly sample 16 frames per video.
Figure 3: We present a comparison on the diving code detection task, where the model need capture each fine-grained movement. The Qwen2.5-VL-7B-Instruct model, relying on passive, single-pass processing of 16 sparsely sampled frames, misses the high-speed contact and made the wrong conclusion. Our DeepSport model, despite having only 7B parameters, active the multi-turn conversation by involking the frame_extraction_tool(35, 59) to retrieve second roun relevant frames. Using this new evidence, it correctly identifies the diving code successfully. This demonstrates the superiority of our active, iterative reasoning paradigm over static models.
4.2 Main Results
We evaluate all models on our 6.7k multi-task, multi-sport test set using an LLM-as-Judge (Claude 4.5 Sonnet [anthropic_2025_introducing]), with scores scaled to a percentage format (0-100).
The comprehensive evaluation results are presented in Table 2. This trend of high performance extends across the diverse task spectrum of our benchmark. When viewed holistically, DeepSport’s well-rounded capabilities become clear. It achieves an overall average score of 40.08, establishing a new state-of-the-art by surpassing the powerful closed-source GPT-5 and the massive Qwen3-VL-235B-A22B-Thinking.
Quantitatively, DeepSport’s superiority is most pronounced in perception-heavy and rule-bound tasks. It secures a dominant lead in Fine-Grained Recognition with a score of 51.09 and Rule & Procedural tasks with 43.82. This substantial margin over generalist models validates our core hypothesis: the active tool-use paradigm effectively captures fleeting visual details missed by passive processors, while agentic RL successfully internalizes complex domain-specific regulations.
In terms of efficiency, while the Qwen3-VL-235B-A22B-Thinking model retains a marginal edge in generative tasks like Assessment & Coaching and Live Commentary—likely leveraging its massive pre-trained knowledge base—DeepSport delivers highly competitive performance despite having only $\sim$3% of the parameter count. This signifies that our distillation pipeline effectively transfers high-level reasoning capabilities to a compact backbone. Moreover, our model achieves the SOTA overall score using only 14.39 frames on average, fewer than the fixed 16 frames used by other models.
Furthermore, the necessity of specialized training is underscored by the comparison with general-domain baselines. DeepSport achieves a score improvement of 23.1 over its own backbone, Qwen2.5-VL-7B, and significantly outperforms the general RL-trained Video-R1 model. This performance gap underscores that general-domain training fails to capture sport-specific nuances, highlighting the critical need for our multi-task specialized model. Figure 3 shows the output comparison between our DeepSport model and the backbone model, Qwen2.5-VL-7B.
4.3 Generalization on General Benchmarks
To investigate whether our sport-specific training compromises general video understanding capabilities, we evaluated DeepSport on two widely used general benchmarks: MLVU [zhou2024mlvu] and Video-MME (Long) [fu2024video]. As shown in Table 3, DeepSport exhibits a slight performance drop compared to its backbone, Qwen2.5-VL-7B-Instruct. This marginal degradation is a common trade-off in domain-specific fine-tuning, which has been discussed in luo2025empirical work. However, notably, DeepSport achieves this competitive performance using significantly fewer frames ($\sim$18 frames) compared to the fixed 32 frames of the backbone. This demonstrates that our active frame extraction mechanism generalizes beyond sports, allowing the model to efficiently locate key information in general long videos while maintaining robust performance superior to previous baselines like VideoChat2 [li2023videochat].
Table 3: Performance on general video understanding benchmarks. DeepSport maintains competitive results with lower average frame usage.
Model MLVU Test Video-MME (Long)
Score Avg. Frames Score Avg. Frames
VideoChat2 35.10 16.00 33.20 16.00
Qwen2.5-VL-7B-Instruct 41.60 32.00 41.90 32.00
DeepSport (Ours) 39.62 18.97 39.98 18.37
4.4 Error Analysis
To investigate the limitations of DeepSport, we conducted a manual inspection of 70 randomly sampled failure cases. We classify errors based on the the first and primary error the model made, categorizing them into six primary types: (1) Tool Grounding Failure (42.9%): The model invoked the tool but retrieved a temporal window that missed the critical event, making the answer impossible. (2) Visual Hallucination (37.1%): The model successfully retrieved the correct frames but misidentified objects, actions, or attributes. (3) Domain Knowledge Gap (11.4%): The model perceived the visual information correctly but lacked the specific sports terminology to describe it. (4) Rule Misapplication (4.3%): The model understood the visual scene and terms but failed in the logical reasoning required to apply the rule. (5) Internal Consistency Error (2.9%): The model’s reasoning contradicted its final answer or violated the output format. (6) Incomplete Exploration (1.4%): The model failed to invoke the frame extraction tool despite the query requiring additional visual evidence.
Figure 4: Error analysis on 70 sampled failure cases. The dominance of Tool Grounding Failure (42.9%) highlights that precise temporal localization remains the primary bottleneck, followed by fine-grained Visual Hallucination (37.1%).
As shown in Figure 4, the dominance of Tool Failure indicates that while our model is capable of reasoning, the primary bottleneck remains accurate temporal localization in long videos. Furthermore, the high rate of Visual Hallucination suggests that future work should focus on enhancing the fine-grained visual encoder resolution.
5 Conclusion
In this paper, we introduced DeepSport, the first end-to-end trained MLLM model designed to unify multi-task video understanding across diverse sports. By synergizing a data distillation pipeline with tool-based agentic reinforcement learning, DeepSport shifts the paradigm from passive frame processing to active, iterative multi-turn reasoning. Our extensive experiments demonstrate that this approach significantly outperforms existing state-of-the-art models, establishing a new benchmark for sport-specific video understanding.
Limitations and Future Work.
Despite these advancements, our work also highlights the inherent limitations of the current sports AI landscape, primarily stemming from data sparsity. While we have curated a comprehensive set of tasks to date, the availability of high-quality, publicly accessible data remains uneven across different sports. For instance, while soccer benefits from abundant commentary and event data, niche sports like fencing or diving often lack such dense, linguistic annotations, making tasks like “fencing commentary generation” currently infeasible. While our framework currently covers a representative range of major sports, numerous less-popular ones remain uncovered.
Second, our error analysis reveals that Tool Grounding Failure remains the primary issue, indicating that while the model effectively decides when to use tools, its ability to pinpoint exact temporal windows needs further refinement. Future work will focus on improving the temporal precision of the retrieval module and developing more comprehensive, multi-granular datasets to pave the way for a truly universal sports intelligence.